OpenStudy (anonymous):

Isosceles triangle ABC is similar to the triangle MNP with vertex angles A and M respectively. If vertex angle A measures 56 degrees, which is the measure of the side angle N? Answer 48 degrees 62 degrees 68 degrees 72 degrees

OpenStudy (anonymous):

62 degrees

OpenStudy (anonymous):

180 - 56 = 124 124/2 = 62

OpenStudy (anonymous):

depending on which angle you are given it could be 68 if A=M then 180-56*2=68

OpenStudy (anonymous):

True, but it states: If vertex angle A measures 56 degrees, ...